# Meta-Reasoner: Dynamic Guidance for Optimized Inference-time Reasoning in Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) increasingly rely on prolonged reasoning chains to solve complex tasks. However, this trial-and-error approach often leads to high computational overhead and error propagation, where early mistakes can derail subsequent steps. To address these issues, we introduce Meta-Reasoner, a framework that dynamically optimizes inference-time reasoning by enabling LLMs to “think about how to think.” Drawing inspiration from human meta-cognition and dual-process theory, Meta-Reasoner operates as a strategic advisor, decoupling high-level guidance from step-by-step generation. It employs contextual multi-armed bandits to iteratively evaluate reasoning progress, and select optimal strategies (e.g., backtrack, clarify ambiguity, restart from scratch, or propose alternative approaches), and reallocates computational resources toward the most promising paths. Our evaluations on mathematical reasoning and puzzles highlight the potential of dynamic reasoning chains to overcome inherent challenges in the LLM reasoning process and also show promise in broader applications, offering a scalable and adaptable solution for reasoning-intensive tasks.
@@
Yuan Sui1, Yufei He1, Tri Cao1, Simeng Han2, Bryan Hooi1 1 National University of Singapore 2 Yale University {yuansui, bhooi}@comp.nus.edu.sg
@@ 1 Introduction
o1-like reasoning chains allow Large Language Models (LLMs) to “think for an extended period” before actually solving a problem. This shows impressive performance on challenging tasks, such as logical problems puzzles Lei et al. (2024); Yao et al. (2023), math questions Patel et al. (2024); Lightman et al. (2023), logical reasoning Han et al. (2024), and science questions Rein et al. (2023), which often pose difficulties for even the most advanced models Gandhi et al. (2024); Sui et al. (2024c); He et al. (2024).
However, the trial-and-error nature of o1-like reasoning often incurs substantial computational overhead Snell et al. (2024); Manvi et al. (2024) and is prone to error propagation where early flaws in a reasoning chain can compound and derail subsequent steps Lei et al. (2024); Yao et al. (2023); Gandhi et al. (2024). While related iterative approaches Gandhi et al. (2024); Li et al. (2025) have explored techniques like partial revision or backtracking, they typically address errors in an ad-hoc manner for a narrow span of reasoning steps and lack a systematic way to assess whether an entire line of reasoning remains viable. Thus, models remain vulnerable to getting “stuck” in less promising reasoning trajectories, continuously expending computational resources on unpromising paths rather than recognizing when a major strategic shift is needed. A critical challenge, therefore, is to enable LLMs to allocate their reasoning budget more effectively, prioritizing promising avenues while adapting—or discarding—ineffective strategies.
Figure 1: High-level comparison of LRM with o1-like chains and meta-reasoner.
@@
To overcome these challenges, we propose Meta-Reasoner, a specialized module that operates alongside the LLM to enhance its reasoning capabilities. The meta-reasoner serves as an “advisor”, dynamically evaluates the reasoning process, offering high-level guidance and strategic redirection when progress stalls. Unlike the LLM, which focuses on more specific stepwise generation, the meta-reasoner focuses on a broader perspective and evaluates the overall progress and strategy of the reasoning process. Meta-Reasoner operates in iterative rounds: where the LLM first generates partial o1-like reasoning chains and provides a “progress report” summarizing its reasoning progress so far. Based on these reports, the meta-reasoner then provides the strategic guidance, such as restarting with a different approach, refining existing ideas, or targeting specific sub-problems. Crucially, the meta-reasoner is designed to not interact with the granular details of the CoT; instead, it focus on the global oversight of the reasoning progress and provides dynamic high-level strategies, preventing the LLM from getting stuck or wasting resources on unproductive lines of inquiry.
Overall, the main contributions of this paper are summarized as follows:
  • We propose a novel meta-reasoning framework that operates as a high-level advisor for LLMs, enabling them to “think about how to think” by dynamically optimizing inference-time reasoning strategies.
  • By decoupling global strategy decisions from low-level chain-of-thought generation, Meta-Reasoner oversees progress through concise “progress reports” rather than micromanaging each reasoning step. This design mitigates error propagation and reduces wasted computation on unproductive paths.
@@
  • We evaluate Meta-Reasoner on challenging mathematical and scientific reasoning benchmarks (e.g., Game of 24, TheoremQA, and SciBench), demonstrating significant improvements in both accuracy and efficiency compared to baselines. Our results show that the proposed framework offers a scalable solution to inference-time reasoning bottlenecks.
@@ 2 Related Works
The introduction of CoT reasoning has revolutionized how LLMs approach problem-solving, allowing them to break tasks into intermediate steps (Lee et al., 2025). The recent LLMs like o1, o3 from OpenAI and Deepseek-v3 from Deepseek have achieved state-of-the-art results in diverse domains using CoT-like reasoning (Manvi et al., 2024; Li et al., 2025; Kudo et al., 2024; Sui et al., 2024b). However, CoT’s sequential dependency limits its robustness, as errors in earlier steps can cascade through the process (Snell et al., 2024) and also when facing complex reasoning tasks (Sui et al., 2024a, c), CoT-like reasoning may stuck in the infinite loop of reasoning (Lee et al., 2025). These issues motivate us to propose Meta-Reasoner to assess and adapt the overall reasoning strategy based on the progress of CoT reasoning. Unlike the LLMs, which focuses on more specific each step generation, the meta-reasoner focus on the border perspective and evaluates the overall progress and strategy of the reasoning process. It can provide a global oversight to avoid LLMs getting stuck or wasting resources on unproductive lines of thoughts.
Backtracking and Error Correction
Addressing cascading errors in multi-step reasoning remains a central challenge. Recent approaches have focused on backtracking and self-verification Yao et al. (2023); Besta et al. (2023); Gandhi et al. (2024). For instance, Weng et al. (2023) have shown that incorporating a self-verification stage—where the model re-checks its conclusions using the very chain of thought it generated—can dramatically boost performance by catching missteps early. Similarly, Ling et al. (2023) propose not only generate multiple candidate reasoning chains but also employs a verifier mechanism to identify and backtrack on erroneous steps. These techniques go beyond post-hoc validation by introducing dynamic strategy adjustments during inference Lightman et al. (2023), thereby reducing the impact of errors propagating through long reasoning chains. Following these useful efforts, we initiate our Meta-Reasoner with the instructions like (1) restart from scratch and propose alternative strategies; (2) backtracking to the point where the error occurred; and (3) continue and provide specific suggestions. The detailed strategy can be found in §4.3.
@@
Meta-cognition in human reasoning involves higher-order processes that allow individuals to monitor, evaluate, and adjust their cognitive strategies (Gao et al., 2024; Yoran et al., 2024). This reflective thinking—often seen as System 2 processes in dual-process theories Havrilla et al. (2024)—is vital for tasks that require careful deliberation and error correction Didolkar et al. (2024). Drawing on these insights, our Meta-Reasoner can be considered analogous to dual-process systems, where LRM for generating CoT steps parallels System 1 and Meta-Reasoner for providing high-level strategic oversight to guide or redirect reasoning when needed serves as System 2. This separation of responsibilities enables the framework to balance efficiency with robust problem-solving, where the LRM handles routine inferences and the meta-reasoner intervenes when high-level strategic adjustments.
@@ 3 Preliminary
A central challenge in complex reasoning tasks is choosing the most effective strategy among multiple valid options. This challenge naturally aligns with the contextual multi-armed bandit (MAB) framework, which is designed to balance the exploration of new strategies with the exploitation of strategies known to perform well.
In this framework, an agent observes a context $x_{t}$ that characterizes the current state of the environment at every time step $t$ and selects an arm $s_{t}$ from a finite set $\mathcal{S}$. Upon selecting arm $s_{t}$, the agent receives a reward $r(s_{t},x_{t})$ that reflects both the chosen arm and the context. The primary objective of MAB is to maximize the cumulative reward over time: $R(T)=\sum_{t=1}^{T}r(s_{t},x_{t})$. A central challenge in the contextual MAB problem is balancing exploration (trying different arms to gather information about their rewards) with exploitation (selecting the arm that has yielded high rewards in similar contexts in the past). This balance ensures that while the agent leverages known profitable actions, it also continues to search for potentially better options. This principle is central to our motivation behind Meta-Reasoner, which aims to automatically select the most efficient strategy to guide the reasoning process during inference time.
A widely used algorithm in the contextual setting is LinUCB Li et al. (2012), which models the expected reward as a linear function of the context. Specifically, for an arm $s$ given context $x_{t}$, the expected reward is modeled as $\mathbb{E}\left[r\left(s,x_{t}\right)\right]\approx x_{t}^{\top}\theta_{s}$, where $\theta_{s}$ is an unknown parameter vector associated with arm $s$. To manage uncertainty in its estimates, LinUCB maintains for each arm an estimate $\hat{\theta}_{s}$ and an associated covariance matrix $A_{s}$. The algorithm then selects the arm according to:
@@
where $c$ is a constant that controls the exploration level. Here, the term $c\sqrt{x_{t}^{\top}A_{s}^{-1}x_{t}}$ serves as a confidence bound on the reward estimate, encouraging the selection of arms with higher uncertainty (i.e., those with less historical data) and thereby facilitating exploration. By incorporating context into the decision-making process, LinUCB allows the agent to adapt its strategy based on the current state, aligning well with our goal of selecting the most efficient reasoning strategy under varying conditions.
@@ 4 Methods
Based on the intuition to allow the LLMs to focus on their computation on more promising lines, we are motivated by two research questions: (1) How can we enable language models to dynamically allocate resources during inference to optimize for reasoning and planning? (2) What architecture allows for effective separation between the reasoning process in LRM and the meta-level guidance of that process in Meta-reasoner? To address them, we propose a new framework, Meta-Reasoner, to equip the LLMs with the capabilities to “thinking about how to think”. It supervises the reasoning process of the LLMs and provides dynamic strategies to enable the LLMs to focus on more promising reasoning paths instead of iterative “trial-and-error”. This framework also addresses limitations of the current sequential generation of the reasoning paths which may get stuck in suboptimal trajectories by balancing “higher-order” thinking.
@@
The meta-reasoning framework operates iteratively as shown in Figure 2. At each round $t$, the reasoning process comprises three steps: (1) CoT generation by the LLM, (2) Progress Reporting to summarize the reasoning progress so far (i.e., this is partly for efficiency, and partly to help the meta-reasoner focus on its main goal of “advising” rather than being distracted by the details in the CoT), and (3) Strategy Generation by the meta-reasoner to optimize subsequent steps. The selection of the strategy is almost exactly corresponds to the well-studied problem of contextual multi-armed bandits learning. Each strategy can be seen as an arm for the bandit, and the reward of each strategy can be evaluated by the progress of LLM reasoning after applying the strategy. We analogy the process of executing and evaluating each strategy as the act of “pulling” each arm. The overall goal of our meta-reasoner is to find the best arm (i.e., strategy with highest cumulative rewards) with as few pulls as possible. The complete algorithm of Meta-Reasoner is appended in Algorithm 1.
@@ 4.1 Chain-of-Thought (CoT) Generation
In the first step, the LLM generates a reasoning step to extend its CoT reasoning based on the user query. Starting with its reasoning history $C_{t-1}$ and the guidance $G_{t-1}$ provided by the meta-reasoner in the previous round, the LRM $M$ produces a new reasoning step $s_{t}$. This step is then appended to the current CoT, forming $C_{t}=C_{t-1}\cup\{s_{t}\}$. This incremental process allows the LRM to iteratively build a structured reasoning path. By keeping track of the full reasoning trajectory at each round, the model creates a coherent foundation for evaluation and further refinement. This process is alike the demonstration in o1-like models, which generate a long-term thinking process. However, the issue of this reasoning is that its more like a process of “trial-and-error”, which may waste some of the inference costs on unnecessary/useless paths. In addition, due to the sequential generation process, it may easily get stuck in suboptimal solutions.
@@ 4.2 Progress Reporting
Once the LRM has updated its CoT, we summarize the reasoning history $C_{t}$ into a concise progress report $P_{t}$. This summary captures the key aspects of the reasoning trajectory, such as how much progress has been made toward the task goal, the consistency of the reasoning, and any significant updates so far. The summarization function $f$ abstracts the detailed CoT into a simpler, more focused representation. This step is designed to be both computationally efficient and informative, ensuring that the meta-reasoner can focus on evaluating high-level progress without being overwhelmed by the granular details of every reasoning step. Even this step is more like an engineering trick, but we find that it may unlock some of the capabilities of LRM to do “higher-order” thinking, we find that with more essential information included in the prompt, the LRM will generally produce more insightful, critical strategies which could be more useful for complex reasoning.
@@
Evidence of error propagation or low confidence in certain subcomponents. Perform targeted verification on critical steps and focus on areas with low confidence.
@@
Table 1: Demonstration: Contextual bandit pair (i.e., diagnosis of the current state and corresponding strategy) for guiding the LLM’s reasoning process. Marked rows are some of the unique strategies generated by Dynamic Contextual Bandits.

### v2

@@ Abstract
Large Language Models (LLMs) increasingly rely on prolonged reasoning chains to solve complex tasks. However, this trial-and-error approach often leads to high computational overhead and error propagation, where early mistakes can derail subsequent steps. To address these issues, we introduce Meta-Reasoner, a framework that dynamically optimizes inference-time reasoning by enabling LLMs to “think about how to think.” Drawing inspiration from human meta-cognition and dual-process theory, Meta-Reasoner operates as a strategic advisor, decoupling high-level guidance from step-by-step generation. It employs contextual multi-armed bandits to iteratively evaluate reasoning progress and select optimal strategies (e.g., backtrack, clarify ambiguity, restart from scratch, or propose alternative approaches), and reallocates computational resources toward the most promising paths. Our evaluations on mathematical reasoning and puzzles highlight the potential of dynamic reasoning chains to overcome inherent challenges in the LLM reasoning process and also show promise in broader applications, offering a scalable and adaptable solution for reasoning-intensive tasks.
@@
Yuan Sui1, Yufei He1, Tri Cao1, Simeng Han2, Yulin Chen1, Bryan Hooi1 1 National University of Singapore 2 Yale University yuan.sui@u.nus.edu, bhooi@comp.nus.edu.sg
@@ 1 Introduction
Recent advances in o1-like and r1-like [^1] O1 (from OpenAI) and R1 (from Deepseek) are two advanced LLM models designed for reasoning. reasoning have enabled large language models (LLMs) to achieve remarkable performance on complex tasks such as mathematics Patel et al. (2024); Lightman et al. (2023), science Rein et al. (2023), and logical puzzles Lei et al. (2024); Yao et al. (2023). By simulating multi-step, human-like deliberation Yao et al. (2024), these methods allow LLMs to decompose problems into smaller subproblems, test hypotheses, reflect on intermediate results, and iteratively refine their solutions. This extended reasoning process enables systematic exploration of ideas, verification of partial conclusions, and progressive improvement before producing a final answer. Such capabilities are particularly valuable in domains demanding rigorous logical reasoning Chenghao Yang (2024).
Despite these advances, o1/r1-like reasoning remains fundamentally challenged by its trial-and-error nature: models generate numerous candidate reasoning paths, discard flawed ones, and gradually converge on solutions. While this flexibility facilitates exploration of diverse strategies, it often incurs substantial computational overhead Snell et al. (2024); Manvi et al. (2024) and is vulnerable to error propagation, where early mistakes accumulate and compromise subsequent steps Lei et al. (2024); Yao et al. (2023); Gandhi et al. (2024). Some iterative methods incorporate partial revision or backtracking Gandhi et al. (2024); Li et al. (2025a), but these approaches tend to be ad-hoc and limited to correcting errors within a narrow reasoning window. Crucially, they lack a systematic mechanism to assess whether an entire reasoning trajectory remains promising or should be abandoned.
As a result, LLMs risk becoming “stuck” on unproductive reasoning paths, wasting valuable computational resources without recognizing when a strategic pivot is necessary. A critical challenge, therefore, is to enable LLMs to manage their reasoning budget more effectively—prioritizing promising directions while adapting or discarding ineffective strategies during inference time.
@@
To address this challenge, we propose Meta-Reasoner, a specialized meta-reasoning module that operates alongside the LLM to enhance its reasoning capabilities. Acting as a high-level advisor, the meta-reasoner dynamically evaluates the reasoning process and provides strategic guidance or redirection when progress stalls. Unlike the LLM, which focuses on detailed stepwise generation, the meta-reasoner maintains a global perspective, assessing overall progress and strategy from a high level. Meta-Reasoner operates in iterative rounds: First, the LLM generates partial chain-of-thought reasoning chains and a concise “progress report” summarizing its current state. Meta-Reasoner then reviews this report and offers high-level feedback - such as restarting reasoning with a different approach, refining existing ideas, or focusing on specific subproblems. This setup allows Meta-Reasoner to concentrate on overall strategy rather than getting involved in the granular details of the LLM’s reasoning. Overall, Meta-Reasoner helps prevent the LLM from getting stuck or spending resources on unproductive lines of inquiry during the inference time.
Overall, the main contributions are as follows:
  • We propose a meta-reasoning module that provides high-level oversight and guidance to the LLM’s reasoning, helping to prevent it from getting stuck on unproductive solution paths during inference.
  • We design a lightweight progress reporting mechanism, enabling efficient communication of the LLM’s reasoning state to the meta-reasoner with minimal overhead.
@@
  • We empirically show that our Meta-Reasoner improves accuracy and efficiency on challenging math and scientific reasoning benchmarks (e.g., Game of 24, TheoremQA, and SciBench) compared to existing strong baselines. Our results highlight the framework’s potential as a new scalable solution to inference-time reasoning.
@@ 2 Related Works
In this section, we first review the challenges of complex reasoning in LLMs and the limitations of chain-of-thought (CoT) methods. We then discuss recent advances in backtracking and self-verification techniques that aim to mitigate these issues. Finally, we draw inspiration from cognitive science, framing our approach within the meta-cognition and dual-process theory to motivate the design of our Meta-Reasoner framework.
@@
The introduction of CoT reasoning has transformed how LLMs tackle complex problems by decomposing tasks into intermediate steps Lee et al. (2025). Recent models such as OpenAI’s o1 and o3, and Deepseek’s r1, have achieved state-of-the-art performance across diverse domains by leveraging CoT-like reasoning Manvi et al. (2024); Li et al. (2025a); Kudo et al. (2024); Sui et al. (2024b). However, the inherent sequential dependency in CoT limits robustness: errors in early steps can propagate and degrade overall performance Snell et al. (2024). Moreover, when confronted with complex reasoning tasks, CoT methods may become trapped in infinite reasoning loops Lee et al. (2025); Sui et al. (2024a, c). To address these challenges, we propose Meta-Reasoner, a meta-reasoner that monitors and adapts the reasoning strategy based on the progress of CoT. By integrating mechanisms such as backtracking and self-verification, Meta-Reasoner provides a holistic perspective that evaluates the overall reasoning trajectory to prevent stagnation and reduce wasted computation on unproductive lines of thought.
Backtracking and Self-Verification
To mitigate the limitations of CoT-like reasoning, recent methods have explored backtracking and self-verification techniques Yao et al. (2023); Besta et al. (2023); Gandhi et al. (2024). For example, Weng et al. (2023) demonstrate that incorporating a self-verification stage—where the model re-examines its conclusions using the generated chain of thought—significantly improves performance by detecting errors early. Similarly, Ling et al. (2023) propose generating multiple candidate reasoning chains alongside a verifier mechanism that identifies and backtracks on erroneous steps. These approaches extend beyond post-hoc validation by enabling dynamic strategy adjustments during inference Lightman et al. (2023), thereby limiting error propagation in lengthy reasoning chains and mitigating infinite reasoning loops. Building on these efforts, our Meta-Reasoner framework employs instructions to (1) restart from scratch with alternative strategies, (2) backtrack to the error point, and (3) continue with targeted suggestions. Further details on this strategy are provided in §4.3.
@@
From a cognitive science perspective, meta-cognition involves higher-order processes that allow individuals to monitor, evaluate, and adjust their cognitive strategies Gao et al. (2024); Yoran et al. (2024). This reflective thinking—often characterized as System 2 in dual-process theories Havrilla et al. (2024)—is vital for tasks requiring careful deliberation and error correction Didolkar et al. (2024). Drawing on these insights, our Meta-Reasoner framework can be viewed as analogous to dual-process systems: the LLM generates CoT steps akin to System 1, while the Meta-Reasoner provides high-level strategic oversight, analogous to System 2, guiding or redirecting reasoning as needed. This separation of responsibilities balances efficiency with robust problem-solving, allowing the LLM to handle routine inferences and the Meta-Reasoner to intervene for strategic adjustments.
@@ 3 Preliminary
In complex reasoning tasks, a key challenge is deciding which strategy to use from multiple valid options. This problem can be naturally framed as a contextual multi-armed bandit (MAB) problem, a well-studied framework for making decisions that balance trying new options (exploration) and using what is already known to work (exploitation).
Imagine an agent faced with several strategies (called arms), and at each step, it observes some information about the current situation (called the context). Based on this context, the agent picks one strategy to apply and then receives feedback (a reward) indicating how well that strategy performed. The goal is to choose strategies over time to maximize the total reward.
Formally, at each time step $t$, the agent observes a context vector $x_{t}$ describing the current state and selects an arm $s_{t}$ from a set of possible strategies $\mathcal{S}$. After choosing $s_{t}$, it receives a reward $r(s_{t},x_{t})$ that depends on both the chosen strategy and the context. The agent aims to maximize the cumulative reward over $T$ steps:
$R(T)=\sum_{t=1}^{T}r(s_{t},x_{t}).$ (1)
A popular algorithm for this problem is LinUCB Li et al. (2012), which assumes the expected reward is approximately a linear function of the context. Specifically, for each arm $s$, there is an unknown parameter vector $\theta_{s}$ such that:
$\mathbb{E}[r(s,x_{t})]\approx x_{t}^{\top}\theta_{s}.$ (2)
LinUCB maintains an estimate $\hat{\theta}_{s}$ of this parameter and a measure of uncertainty about the estimate. At each step, it selects the arm that maximizes a combination of the estimated reward and an uncertainty bonus:
@@
where $A_{s}$ is a matrix capturing past observations for arm $s$, and $c$ controls how much the agent favors exploring uncertain options. The second term encourages trying arms that might perform well but have been less explored.
By using context to guide its choices, LinUCB adapts its strategy selection to the current situation, which aligns with our goal in Meta-Reasoner: to dynamically choose the most effective reasoning strategy during inference time.
@@ 4 Methods
Motivated by the intuition that LLMs should concentrate their computational efforts on more promising reasoning paths during inference time, we explore two key research questions in this paper: (1) How can language models dynamically allocate resources during inference to optimize reasoning and planning?; (2) What architectural design enables an effective separation between the reasoning process within the LLM and the meta-level guidance that oversees it?
To address these questions, we propose a novel framework, Meta-Reasoner, which endows LLMs with the ability to “think about how to think”. Our framework supervises the reasoning process and dynamically guides the model to focus on more promising reasoning trajectories during inference time. Furthermore, Meta-Reasoner mitigates the limitations of conventional sequential reasoning, which may get stuck in suboptimal paths. We propose a “high-order" reasoning mechanism to balance exploration and exploitation using a Multi-Armed Bandit (MAB) algorithm.
@@
The meta-reasoning framework operates iteratively as illustrated in Figure 1. At each round $t$, the reasoning process comprises three steps: (1) CoT generation by the LLM, (2) Progress Reporting to summarize the reasoning progress so far (i.e., this is partly for efficiency, and partly to help the meta-reasoner focus on its main goal of “advising” rather than being distracted by the details in the CoT), and (3) Strategy Generation by the meta-reasoner to optimize subsequent steps. The selection of the strategy is almost exactly corresponds to the well-studied problem of contextual multi-armed bandits as illustrated in §3. Each strategy can be seen as an arm for the bandit, and the reward of each strategy can be evaluated by the progress of LLM reasoning after applying the strategy. We analogy the process of executing and evaluating each strategy as the act of “pulling” each arm. The overall goal of our meta-reasoner is to find the best arm (i.e., strategy with highest cumulative rewards) with as few pulls as possible. The complete algorithm of Meta-Reasoner is appended in Algorithm 1.
@@ 4.1 Chain-of-Thought (CoT) Generation
In the first step, the LLM generates a reasoning step to extend its CoT reasoning based on the user query. Starting from its reasoning history $C_{t-1}$ and the guidance $G_{t-1}$ provided by the meta-reasoner in the previous round, the LLM $M$ produces a new reasoning step $s_{t}$. This step is appended to the current CoT, forming $C_{t}=C_{t-1}\cup\{s_{t}\}$. By maintaining the full reasoning trajectory at each round, the model establishes a coherent foundation for evaluation and further refinement. This approach resembles the long-term reasoning demonstrated in models like o1, which generate extended CoTs. However, this reasoning process often resembles “trial-and-error”, potentially incurring unnecessary inference costs on unproductive paths. Moreover, due to its sequential nature, the process is prone to becoming trapped in suboptimal solutions.
@@ 4.2 Progress Reporting
After updating the CoT, we summarize the reasoning history $C_{t}$ into a concise progress report $P_{t}$. This summary captures key aspects of the reasoning trajectory, including progress toward the task goal, reasoning consistency, and significant updates thus far. The summarization function $f$ abstracts the detailed CoT into a simpler, more focused representation. This step is designed to be both computationally efficient and informative, enabling the meta-reasoner to evaluate high-level progress without being overwhelmed by the granular details of every reasoning step. While this may be viewed as an engineering heuristic, we observe that including essential information in the prompt unlocks the LLM’s capacity for “higher-order” thinking. Consequently, the LLM tends to generate more insightful and critical strategies, which are particularly valuable for complex reasoning tasks.
@@
Evidence of error propagation or low confidence in certain sub-components. Perform targeted verification on critical steps and focus on areas with low confidence.
@@
Table 1: Demonstration: Contextual bandit pair (i.e., diagnosis of the progress report (context) and the corresponding strategy (bandit)) for guiding the LLM’s reasoning process. Marked rows are some of the unique strategies generated by Dynamic Contextual Bandits.